Operation Summary for October 16, 2009 - October 19, 2009 |
The Integrated Luminosity for the period from 10/12/09 to 10/19/09 was 55.11 pb-1. Friday Oct. 16 The day shift began with Tevatron (TeV) personnel investigating the sector F1 quench, with Operations monitoring the Recycler stash (146.6E10), with the Antiproton Source stacking (26.6E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the NuMI and MiniBooNE experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments waiting for the next store. During the day shift, TeV personnel worked to understand the quench at F17L. CDF and D0 accessed their collision halls. TeV personnel conducted beam studies. Operations began the TeV turn-on by noon. MI and Switchyard (SY) personnel successfully transported slow-spill beam to the SY dump. MECAR developed a memory problem that required a reboot and time to re-learn all of the MI ramp resets; ~10E10 antiprotons were lost out of the Recycler stash. During the evening shift, TeV personnel concluded their beam studies. Trouble with cryo system CAMAC cards at F0 and a yard breaker trip at sector E0 delayed TeV orbit checks. Operations put the TeV into HEP shot setup at 8:45 PM and established store 7275 with an average initial luminosity (AIL) of 240.87E30 at 10:30 PM. Saturday Oct. 17 The midnight shift began with Operations monitoring store 7275 (177.61E30) and the stash (81.0E10), with Pbar stacking (13.2E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, the Booster beam efficiency dropped towards the end of shift for no obvious reason. During the day shift, Operations terminated store 7275 at 3 PM. Operations put the TeV into shot setup and established store 7276 at 4:30 PM with an AIL of 268.64E30. During the evening shift, Operations reported problems with MI-8 transfer line beam permit trips; efforts to fix the beam position problems and the frequency of the trips were unsuccessful, though the trips grew less frequent throughout the shift. Sunday Oct. 18 The midnight shift began with Operations monitoring store 7276 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam. During the midnight shift, Operations monitored and tuned, for the majority of the shift, on the Linac and Booster machines to resolve beam position problems; progress was made towards the end of the shift. The Pelletron autotune program required a restore of all its parameters. During the day shift, Operations terminated store 7276 at 8:45 AM. Operations put the TeV into shot setup and established store 7277 at 10:30 AM with an AIL of 272.9E30. During the evening shift, the TeV sector F3 air compressor tripped off, but was reset before the pressure dropped low enough to threaten the store. Monday Oct. 19 The midnight shift began with Operations monitoring store 7277 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam. During the midnight shift, Operations terminated store 7277 at 2:30 AM. Operations put the TeV into shot setup, but the TeV quenched at B0 during the low beta magnet squeeze. Efforts to understand and recover from the quench followed with an investigation of the beam conditions and a possible problem with a quadrupole (B0Q2) power supply. The Plans for Monday The plans for today are to stack, stash, and store. The Run Coordinator will allow Booster and Pbar studies. |
|
More Information Comments and Suggestions |
| last modified 10/19/2009 email Fermilab |
| Security, Privacy, Legal | |